# Observation of an Accreting Planetary-Mass Companion with Signs of Disk–Disk Interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY

### v1

@@ Acknowledgements.
The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1288. OB, IS are funded by the Centre National d’Etudes Spatiales (CNES) through the APR program. This research received funding from the program ANR-22-EXOR-0001 Origins of the Institut National des Sciences de l’Univers, CNRS. TJH acknowledges a Dorothy Hodgkin Fellowship, UKRI guaranteed funding for a Horizon Europe ERC consolidator grant (EP/Y024710/1) and UKRI/STFC grant ST/X000931/1. This project is co-funded by the European Union (ERC, SUL4LIFE, grant agreement No101096293). AF thanks project PID2022-137980NB-I00 funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe”.References
- Allard et al. (2001) Allard, F., Hauschildt, P. H., Alexander, D. R., Tamanai, A., & Schweitzer, A. 2001, ApJ, 556, 357
- Allard et al. (2011) Allard, F., Homeier, D., & Freytag, B. 2011, in Astronomical Society of the Pacific Conference Series, Vol. 448, 16th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, ed. C. Johns-Krull, M. K. Browning, & A. A. West, 91
- Allard et al. (2012) Allard, F., Homeier, D., & Freytag, B. 2012, Philosophical Transactions of the Royal Society of London Series A, 370, 2765
- Alves & Bouy (2012) Alves, J. & Bouy, H. 2012, Astronomy & Astrophysics, 547
- Aoyama et al. (2021) Aoyama, Y., Marleau, G.-D., Ikoma, M., & Mordasini, C. 2021, The Astrophysical Journal, 917, L30, publisher: IOP ADS Bibcode: 2021ApJ…917L..30A
- Aru et al. (2024) Aru, M. L., Maucó, K., Manara, C. F., et al. 2024, A&A, 692, A137
- Aru et al. (2024) Aru, M.-L., Mauco, K., Manara, C. F., et al. 2024, Astronomy & Astrophysics, 687, A93, arXiv:2403.12604 [astro-ph]
- Baraffe et al. (2003) Baraffe, I., Chabrier, G., Barman, T. S., Allard, F., & Hauschildt, P. H. 2003, Astronomy & Astrophysics, 402, 701
- Baraffe et al. (2015) Baraffe, I., Homeier, D., Allard, F., & Chabrier, G. 2015, A&A, 577, A42
- Berné et al. (2022) Berné, O., Habart, E., Peeters, E., et al. 2022, Publications of the Astronomical Society of the Pacific, 134, 054301
- Boccaletti et al. (2020) Boccaletti, A., Di Folco, E., Pantin, E., et al. 2020, Astronomy & Astrophysics, 637, L5
- Boccaletti et al. (2024) Boccaletti, A., Mâlin, M., Baudoz, P., et al. 2024, A&A, 686, A33
- Bowler et al. (2014) Bowler, B. P., Liu, M. C., Kraus, A. L., & Mann, A. W. 2014, ApJ, 784, 65
- Bowler et al. (2011) Bowler, B. P., Liu, M. C., Kraus, A. L., Mann, A. W., & Ireland, M. J. 2011, The Astrophysical Journal, 743, 148
- Cabrit et al. (2006) Cabrit, S., Pety, J., Pesenti, N., & Dougados, C. 2006, A&A, 452, 897
- Carter et al. (2023) Carter, A. L., Hinkley, S., Kammerer, J., et al. 2023, The Astrophysical Journal Letters, 951, L20, publisher: The American Astronomical Society
- Chown et al. (2024) Chown, R., Sidhu, A., Peeters, E., et al. 2024, A&A, 685, A75
- Clarke & Pringle (1993) Clarke, C. J. & Pringle, J. E. 1993, MNRAS, 261, 190
- Currie et al. (2013) Currie, T., Daemgen, S., Debes, J., et al. 2013, The Astrophysical Journal Letters, 780, L30
- Currie et al. (2022) Currie, T., Lawson, K., Schneider, G., et al. 2022, Nature Astronomy, 6, 751, arXiv:2204.00633 [astro-ph]
- Eriksson et al. (2020) Eriksson, S. C., Asensio Torres, R., Janson, M., et al. 2020, Astronomy & Astrophysics, 638, L6
- Escalante (1991) Escalante, V. 1991, Publications of the Astronomical Society of the Pacific, 103, 838
- Fang et al. (2009) Fang, M., van Boekel, R., Wang, W., et al. 2009, A&A, 504, 461
- Flores-Rivera et al. (2023) Flores-Rivera, L., Flock, M., Kurtovic, N. T., et al. 2023, Astronomy & Astrophysics, 670, A126
- Ginski et al. (2025) Ginski, C., Pinilla, P., Benisty, M., et al. 2025, arXiv e-prints, arXiv:2506.05892
- Goicoechea et al. (2024) Goicoechea, J. R., Le Bourlot, J., Black, J. H., et al. 2024, A&A, 689, L4
- Habart et al. (2024) Habart, E., Peeters, E., Berné, O., et al. 2024, Astronomy & Astrophysics, 685, A73, arXiv:2308.16732 [astro-ph]
- Haffert et al. (2019) Haffert, S. Y., Bohn, A. J., de Boer, J., et al. 2019, Nature Astronomy, 3, 749
- Harris et al. (2012) Harris, R. J., Andrews, S. M., Wilner, D. J., & Kraus, A. L. 2012, ApJ, 751, 115
- Hartigan et al. (1995) Hartigan, P., Edwards, S., & Ghandour, L. 1995, ApJ, 452, 736
- Hashimoto et al. (2020) Hashimoto, J., Aoyama, Y., Konishi, M., et al. 2020, The Astronomical Journal, 159, 222
- Haworth et al. (2023) Haworth, T. J., Reiter, M., O’Dell, C. R., et al. 2023, Monthly Notices of the Royal Astronomical Society, 525, 4129
- Hillenbrand (1997) Hillenbrand, L. A. 1997, AJ, 113, 1733
- Husser et al. (2013) Husser, T. O., Wende-von Berg, S., Dreizler, S., et al. 2013, A&A, 553, A6
- Jayawardhana et al. (2003) Jayawardhana, R., Mohanty, S., & Basri, G. 2003, The Astrophysical Journal, 592, 282, arXiv:astro-ph/0303565
- Joergens et al. (2013) Joergens, V., Bonnefoy, M., Liu, Y., et al. 2013, Astronomy & Astrophysics, 558, L7, arXiv:1310.1936 [astro-ph]
- Kirkpatrick et al. (2006) Kirkpatrick, J. D., Barman, T. S., Burgasser, A. J., et al. 2006, The Astrophysical Journal, 639, 1120, publisher: IOP Publishing
- Kounkel et al. (2018) Kounkel, M., Covey, K., Suárez, G., et al. 2018, The Astronomical Journal, 156, 84, publisher: The American Astronomical Society
- Kratter & Lodato (2016) Kratter, K. & Lodato, G. 2016, ARA&A, 54, 271
- Lagrange et al. (2025) Lagrange, A.-M., Wilkinson, C., Mâlin, M., et al. 2025, Evidence for a sub-jovian planet in the young TWA7 disk, arXiv:2502.15081 [astro-ph]
- Luhman et al. (2006) Luhman, K. L., Wilson, J. C., Brandner, W., et al. 2006, ApJ, 649, 894
- Magazzu et al. (1993) Magazzu, A., Martin, E. L., & Rebolo, R. 1993, ApJ, 404, L17
- Manjavacas et al. (2020) Manjavacas, E., Lodieu, N., Béjar, V. J. S., et al. 2020, Monthly Notices of the Royal Astronomical Society, 491, 5925
- Matthews et al. (2024) Matthews, E. C., Carter, A. L., Pathak, P., et al. 2024, Nature, 633, 789
- Maíz Apellániz et al. (2022) Maíz Apellániz, J., Barbá, R. H., Fernández Aranda, R., et al. 2022, Astronomy & Astrophysics, 657, A131
- Mohanty et al. (2005) Mohanty, S., Jayawardhana, R., & Basri, G. 2005, The Astrophysical Journal, 626, 498
- Mullally et al. (2024) Mullally, S. E., Debes, J., Cracraft, M., et al. 2024, The Astrophysical Journal Letters, 962, L32
- Muzerolle et al. (2000) Muzerolle, J., Briceño, C., Calvet, N., et al. 2000, The Astrophysical Journal, 545, L141
- Natta et al. (2004) Natta, A., Testi, L., Muzerolle, J., et al. 2004, Astronomy and Astrophysics, 424, 603, aDS Bibcode: 2004A&A…424..603N
- Phan-Bao et al. (2017) Phan-Bao, N., Bessell, M. S., Nguyen-Thanh, D., et al. 2017, A&A, 600, A19
- Phillips et al. (2020) Phillips, M. W., Tremblin, P., Baraffe, I., et al. 2020, Astronomy & Astrophysics, 637, A38, arXiv:2003.13717 [astro-ph]
- Piqueras et al. (2017) Piqueras, L., Conseil, S., Shepherd, M., et al. 2017, MPDAF - A Python package for the analysis of VLT/MUSE data, arXiv:1710.03554 [astro-ph]
- Planet formation environments collaboration (2025) Planet formation environments collaboration, n. 2025, The past, present and future of observations of externally irradiated disks, arXiv:2502.12255 [astro-ph]
- Reiter & Parker (2022) Reiter, M. & Parker, R. J. 2022, European Physical Journal Plus, 137, 1071
- Ricci et al. (2008) Ricci, L., Robberto, M., & Soderblom, D. R. 2008, The Astronomical Journal, 136, 2136
- Rogers et al. (2024) Rogers, C., De Marchi, G., & Brandl, B. 2024, Astronomy & Astrophysics, 684, L8
- Santamaría-Miranda et al. (2017) Santamaría-Miranda, A., Cáceres, C., Schreiber, M. R., et al. 2017, Monthly Notices of the Royal Astronomical Society, 475, 2994
- Schlaufman (2018) Schlaufman, K. C. 2018, The Astrophysical Journal, 853, 37
- Schmidt et al. (2008) Schmidt, T., Neuhäuser, R., Seifahrt, A., et al. 2008, Astronomy & Astrophysics, 491, 311
- Solf & Boehm (1993) Solf, J. & Boehm, K. H. 1993, ApJ, 410, L31
- White & Basri (2003) White, R. & Basri, G. 2003, The Astrophysical Journal, 582, 1109, arXiv:astro-ph/0209164
- Zeidler (2019) Zeidler, P. 2019, MUSEpack
- Zeidler et al. (2019) Zeidler, P., Nota, A., Sabbi, E., et al. 2019, The Astronomical Journal, 158, 201
- Zhou et al. (2014) Zhou, Y., Herczeg, G. J., Kraus, A. L., Metchev, S., & Cruz, K. L. 2014, The Astrophysical Journal Letters, 783, L17

### v2

@@ Acknowledgements.
The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1288. OB, IS are funded by the Centre National d’Etudes Spatiales (CNES) through the APR program. This research received funding from the program ANR-22-EXOR-0001 Origins of the Institut National des Sciences de l’Univers, CNRS. TJH acknowledges a Dorothy Hodgkin Fellowship, UKRI guaranteed funding for a Horizon Europe ERC consolidator grant (EP/Y024710/1) and UKRI/STFC grant ST/X000931/1. This project is co-funded by the European Union (ERC, SUL4LIFE, grant agreement No101096293). AF thanks project PID2022-137980NB-I00 funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 and by ”ERDF A way of making Europe”. The authors acknowledge G. D. Marleau and Y. Aoyama for useful discussions on the conversion of H recombination lines intensities into accretion rates.References
- Model Atmospheres From Very Low Mass Stars to Brown Dwarfs. In 16th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, C. Johns-Krull, M. K. Browning, and A. A. West (Eds.), ASP, Vol. 448, pp. 91. External Links: Document, 1011.5405, ADS entry Cited by: Appendix H.
- Models of very-low-mass stars, brown dwarfs and exoplanets. Philos. Trans. R. Soc. of London Series A 370 (1968), pp. 2765–2777. External Links: Document, 1112.3591, ADS entry Cited by: Appendix H.
- The Limiting Effects of Dust in Brown Dwarf Model Atmospheres. ApJ 556 (1), pp. 357–372. External Links: Document, astro-ph/0104256, ADS entry Cited by: Appendix H, §3.1.
- Orion revisited. i. the massive cluster in front of the orion nebula cluster. A&A 547 (547) (English). External Links: Document, ISSN 0004-6361 Cited by: §2.
- Comparison of Planetary Hα-emission Models: A New Correlation with Accretion Luminosity. ApJ 917, pp. L30. External Links: ISSN 0004-637X, Link, Document Cited by: Appendix G.
- A tell-tale tracer for externally irradiated protoplanetary disks: Comparing the [C I] 8727 Å line and ALMA observations in proplyds. 692, pp. A137. External Links: Document, 2410.21018, ADS entry Cited by: §E.2.
- Kaleidoscope of irradiated disks: MUSE observations of proplyds in the Orion Nebula Cluster. I. Sample presentation and ionization front sizes. A&A 687, pp. A93 (en). Note: Comment: Submitted to A&A on December 18, 2023. Accepted on April 2, 2024 External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §E.2.
- Evolutionary models for cool brown dwarfs and extrasolar giant planets. The case of HD 209458. A&A 402 (2), pp. 701–712 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: Figure 6, Appendix D, §3.1.
- New evolutionary models for pre-main sequence and main sequence low-mass stars down to the hydrogen-burning limit. A&A 577, pp. A42. External Links: Document, 1503.04107, ADS entry Cited by: Figure 9.
- PDRs4All: A JWST Early Release Science Program on Radiative Feedback from Massive Stars. Publ. Astron. Soc. Pac. 134 (1035), pp. 054301 (en). External Links: ISSN 0004-6280, 1538-3873, Link, Document Cited by: Appendix A, §1.
- Possible evidence of ongoing planet formation in AB Aurigae: A showcase of the SPHERE/ALMA synergy. A&A 637, pp. L5 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §3.3.
- Imaging detection of the inner dust belt and the four exoplanets in the HR 8799 system with JWST’s MIRI coronagraph. A&A 686, pp. A33. External Links: Document, 2310.13414, ADS entry Cited by: §1.
- A disk around the planetary-mass companion gsc 06214-00210 b: clues about the formation of gas giants on wide orbits*. ApJ 743 (2), pp. 148. External Links: Document, Link Cited by: §3.2.
- Spectroscopic Confirmation of Young Planetary-mass Companions on Wide Orbits. ApJ 784 (1), pp. 65. External Links: Document, 1401.7668, ADS entry Cited by: §E.2, §1.
- Tidal stripping and disk kinematics in the RW Aurigae system. 452 (3), pp. 897–906. External Links: Document, ADS entry Cited by: §3.3.
- The JWST Early Release Science Program for Direct Observations of Exoplanetary Systems I: High-contrast Imaging of the Exoplanet HIP 65426 b from 2 to 16 μm. ApJ 951 (1), pp. L20 (en). External Links: ISSN 2041-8205, Link, Document Cited by: §1.
- Accretion disc response to a stellar fly-by. 261 (1), pp. 190–202. External Links: Document, ADS entry Cited by: §3.3.
- DIRECT imaging and spectroscopy of a candidate companion below/near the deuterium-burning limit in the young binary star system, roxs 42b. ApJ 780 (2), pp. L30. External Links: Document, Link Cited by: §1.
- Images of Embedded Jovian Planet Formation At A Wide Separation Around AB Aurigae. Nat. Astron. 6 (6), pp. 751–759. Note: Comment: Author’s personal version: 19 pages, 5 Figures, 1 Table; 32 Supplementary pages, 18 Supplementary Figures, 1 Supplementary Table; Accepted for Publication in Nature Astronomy. Published version: https://www.nature.com/articles/s41550-022-01634-x External Links: ISSN 2397-3366, Link, Document Cited by: §1.
- Strong H α emission and signs of accretion in a circumbinary planetary mass companion from MUSE. A&A 638, pp. L6 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §E.1, §E.2, §3.2.
- PHOTODISSOCIATION regions. 103 (666), pp. 838. External Links: Document, Link Cited by: §E.2.
- Star and protoplanetary disk properties in Orion’s suburbs. A&A 504 (2), pp. 461–489. External Links: Document, 0907.2380, ADS entry Cited by: §3.2.
- Forbidden emission lines in protostellar outflows and jets with muse. 670, pp. A126. Cited by: §E.2.
- Disk Evolution Study Through Imaging of Nearby Young Stars (DESTINYS): Evidence of planet–disk interaction in the 2MASSJ16120668-3010270 system. A&A 699, pp. A237. External Links: Document, 2506.05892, ADS entry Cited by: §3.3.
- PDRs4All: X. ALMA and JWST detection of neutral carbon in the externally irradiated disk d203-506: Undepleted gas-phase carbon. 689, pp. L4. External Links: Document, 2408.06279, ADS entry Cited by: §E.2.
- PDRs4All II: JWST’s NIR and MIR imaging view of the Orion Nebula. A&A 685, pp. A73 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: Appendix A, Figure 1, §1, §2, §3.
- Two accreting protoplanets around the young star PDS 70. 3, pp. 749–754. External Links: Document, 1906.01486, ADS entry Cited by: §3.3.
- A Resolved Census of Millimeter Emission from Taurus Multiple Star Systems. 751 (2), pp. 115. External Links: Document, 1203.6353, ADS entry Cited by: §3.3.
- Disk Accretion and Mass Loss from Young Stars. 452, pp. 736. External Links: Document, ADS entry Cited by: §E.2.
- Accretion properties of pds 70b with muse*. AJ 159 (5), pp. 222. External Links: ISSN 1538-3881, Link, Document Cited by: §3.2, §3.3.
- The VLT MUSE NFM view of outflows and externally photoevaporating discs near the orion bar★. MNRAS 525 (3), pp. 4129–4142 (en). External Links: ISSN 0035-8711, 1365-2966, Link, Document Cited by: §E.2.
- On the Stellar Population and Star-Forming History of the Orion Nebula Cluster. AJ 113, pp. 1733–1768. External Links: Document, ADS entry Cited by: §2.
- A new extensive library of PHOENIX stellar atmospheres and synthetic spectra. 553, pp. A6. External Links: Document, 1303.5632, ADS entry Cited by: Appendix F.
- Evidence for a T Tauri Phase in Young Brown Dwarfs. ApJ 592 (1), pp. 282–287 (en). Note: Comment: to appear in The Astrophysical Journal External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §E.2.
- OTS44: Disk and accretion at the planetary border. A&A 558, pp. L7 (en). Note: Comment: Accepted for publication in Astronomy and Astrophysics Letter External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §E.2, §3.2.
- Discovery of a Very Young Field L Dwarf, 2MASS J01415823–4633574*. ApJ 639 (2), pp. 1120 (en). External Links: ISSN 0004-637X, Link, Document Cited by: §3.2.
- The APOGEE-2 Survey of the Orion Star-forming Complex. II. Six-dimensional Structure. AJ 156 (3), pp. 84 (en). External Links: ISSN 1538-3881, Link, Document Cited by: §2.
- Gravitational Instabilities in Circumstellar Disks. ARA&A 54, pp. 271–311. External Links: Document, 1603.01280, ADS entry Cited by: §1.
- Evidence for a sub-Jovian planet in the young TWA 7 disk. Nature 642 (8069), pp. 905–908. External Links: Document, 2502.15081, ADS entry Cited by: §1.
- Discovery of a Young Substellar Companion in Chamaeleon. ApJ 649 (2), pp. 894–899. External Links: Document, astro-ph/0609187, ADS entry Cited by: §1.
- A Spectroscopic Test for Substellar Objects. ApJ 404, pp. L17. External Links: Document, ADS entry Cited by: §E.2.
- The Villafranca catalog of Galactic OB groups: II. From Gaia DR2 to EDR3 and ten new systems with O stars. A&A 657, pp. A131 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §2.
- Spectral library of age-benchmark low-mass stars and brown dwarfs. MNRAS 491 (4), pp. 5925–5950. External Links: ISSN 0035-8711, Link, Document Cited by: Figure 10, Appendix F, §3.2.
- A temperate super-Jupiter imaged with JWST in the mid-infrared. Nature 633 (8031), pp. 789–792. External Links: Document, 2503.01599, ADS entry Cited by: §1.
- The t tauri phase down to nearly planetary masses: echelle spectra of 82 very low mass stars and brown dwarfs. ApJA&AApJApJNat. Astron.A&AA&AA&AApJMNRASPubl. Astron. Soc. Pac.A&AMNRASA&AEur. Phys. J. 626 (1), pp. 498. External Links: Document, Link Cited by: §3.2.
- JWST directly images giant planet candidates around two metal-polluted white dwarf stars. ApJ 962 (2), pp. L32. External Links: Document, Link Cited by: §1.
- Detection of Disk Accretion at the Substellar Limit. ApJ 545 (2), pp. L141–L144 (en). External Links: ISSN 0004637X, Link, Document Cited by: §E.2.
- Accretion in brown dwarfs: An infrared view. A&A 424, pp. 603–612. External Links: ISSN 0004-6361, Link, Document Cited by: §E.2.
- Detection of lithium in nearby young late-M dwarfs. A&A 600, pp. A19. External Links: Document, 1612.09370, ADS entry Cited by: §E.2.
- A new set of atmosphere and evolution models for cool T-Y brown dwarfs and giant exoplanets. A&A 637, pp. A38 (en). Note: Comment: 21 pages, 18 figures, 1 table. Accepted for publication in A&A. Models available at http://opendata.erc-atmo.eu and http://perso.ens-lyon.fr/isabelle.baraffe/ATMO2020/ External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: Appendix H.
- MPDAF - A Python package for the analysis of VLT/MUSE data. arXiv. Note: Comment: to be published in ADASS XXVI (held October 16-20, 2016) proceedings. See http://www.adass2016.inaf.it/images/posters/piqueras.pdf for the poster External Links: Link, Document Cited by: §E.2, Appendix G.
- The past, present and future of observations of externally irradiated disks. Open J. Astrophys. 8. External Links: Document Cited by: §1.
- Dynamics of young stellar clusters as planet-forming environments. 137 (9), pp. 1071. External Links: Document, 2209.03889, ADS entry Cited by: §1.
- THE HUBBLE SPACE TELESCOPE /ADVANCED CAMERA FOR SURVEYS ATLAS OF PROTOPLANETARY DISKS IN THE GREAT ORION NEBULA. AJ 136 (5), pp. 2136–2151 (en). External Links: ISSN 0004-6256, 1538-3881, Link, Document Cited by: Appendix A, Figure 1, §2, §3.
- Accretion signatures in the x-shooter spectrum of the substellar companion to sr12. MNRAS 475 (3), pp. 2994–3003. External Links: ISSN 0035-8711, Document, Link, https://academic.oup.com/mnras/article-pdf/475/3/2994/23799375/stx3325.pdf Cited by: §3.2.
- Evidence of an upper bound on the masses of planets and its implications for giant planet formation. ApJ 853 (1), pp. 37. External Links: Document, Link Cited by: §1.
- Direct evidence of a sub-stellar companion around ct chamaeleontis. A&A 491 (1), pp. 311–320. Cited by: §1.
- High-Resolution Long-Slit Spectral Imaging of the Mass Outflows in the Immediate Vicinity of DG Tauri. 410, pp. L31. External Links: Document, ADS entry Cited by: §E.2.
- Dust Grain-Size Distributions and Extinction in the Milky Way, Large Magellanic Cloud, and Small Magellanic Cloud. ApJ 548 (1), pp. 296–309. External Links: Document, astro-ph/0008146, ADS entry Cited by: §3.1.
- Very Low Mass Stars and Brown Dwarfs in Taurus-Auriga. ApJ 582 (2), pp. 1109–1122 (en). Note: Comment: 27 pages, plus 8 figures, accepted for publication in ApJ External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §E.2.
- The Young Massive Star Cluster Westerlund 2 Observed with MUSE. II. MUSEpack—A Python Package to Analyze the Kinematics of Young Star Clusters. AJ 158 (5), pp. 201. External Links: Document, ISSN 1538-3881 Cited by: Appendix A.
- MUSEpack External Links: Document, ADS entry Cited by: Appendix A.
- ACCRETION onto planetary mass companions of low-mass young stars. ApJ 783 (1), pp. L17. External Links: Document, Link Cited by: §3.2.